OpenStudy (anonymous):

72) If two points in a number line have coordinates 1 & 7, find the coordinate of a point on the line which is twice as far from 1 and 7. a) 2 b) 3 c) 4 d) 5

OpenStudy (anonymous):

help

OpenStudy (anonymous):

maybe it should say: If two points in a number line have coordinates 1 & 7, find the coordinate of a point on the line which is twice as far from 1 than 7 ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what is the answer? please...

OpenStudy (anonymous):

let the number be x so 2(x-1)=2(7-x)..giving x=4

OpenStudy (anonymous):

so the answer is 4??

OpenStudy (anonymous):

yes should be..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last line of the question is confusing me though